OpenStudy (anonymous):

There are 15 students on the Prom Committee and 20 students on the Yearbook Committee. If 10 students are on both committees, how many students are on only one committee?

OpenStudy (anonymous):

no of students who r only in one commitee= 15+20-10=25

OpenStudy (anonymous):

(15-10)+(20-10)=15 If you think about it, 5 of the prom committee and 10 yearbook committee are not on both committees.

OpenStudy (anonymous):

i dnt agree @monty5658